# UGround: Towards Unified Visual Grounding with Unrolled Transformers

### v1

@@ 4.1 Policy-Prompted Masking
Algorithm (1) outlines one forward pass of the PPM. As the transformer is composed of stacked layers, selecting a layer $\ell^{*}$ effectively skips the subsequent $L$-$\ell^{*}$ layers, thereby creating a direct connection to SAM and forming a skip-connection–like paradigm that mitigates the dilution of semantic information by noise in intermediate layers. Across multiple forward passes ($1$ to $\mathcal{T}$), this mechanism virtually connects all layers to SAM, with only one path active per pass in a dropout-like fashion. The current mechanism offers the following benefits: 1) Analogous to Monte Carlo Dropout, a layer $\ell^{*}$ is randomly selected in each forward pass, and multiple $\mathcal{T}$ stochastic samples facilitate an approximate Bayesian uncertainty estimation; 2) Sampling across different paths alleviates the model’s over-reliance on any single trajectory; 3) The sampling alters the model’s connectivity with SAM, virtually forming “ensemble” subnetworks. All together, these contribute to the model’s robustness.
Mask as Prompt. Compared with <SEG> tokens, the similarity map provides more explicit spatial cues, as analyzed in Sec. 3.2. We use the similarity map as a logit mask to prompt SAM. Accordingly, we reformulate Eq. (3) as $\hat{\mathbf{M}}=\;\mathcal{G}_{\mathcal{V}}^{dec}(\mathbf{f},\bm{h}_{seg},\mathcal{M}).$ The similarity map is continuously differentiable, allowing its gradients to be backpropagated through the SAM module for implicit optimization. To further guide the model to focus on target regions, we supervise the similarity map $\mathcal{M}$ by imposing a constraint (e.g., using BEC or Dice loss) against a soft ground-truth mask $M_{\sigma}$, providing explicit guidance on where the model should “attend”.
@@ 4.2 Training Objectives
5 Experiment
5.1 Experimental Setting
Implementation Details.
We observe that the center scaling and cropping operation in CLIP may crop out target objects. To address this issue and to align the similarity map with SAM’s input, we adopt SAM’s transformation strategy, scaling the longer side of the image to 336 and padding the image for the CLIP input $\times336$. For results on ReasonSeg, UGround-7B-LLaVA1.5 is trained on 1 NVIDIA A100 GPU 80GB for about 2 days, UGround-13B-LLaVA1.5 for about 13 hours. The training relies solely on ReasonSeg and a referring segmentation dataset ($\sim$10k images). We use LoRA (Hu et al., 2022) for efficient fine-tuning, using lora_r=8lora\_r=8 for 7B and 64 for 13B. We use the AdamW (Loshchilov & Hutter, 2019) optimizer with an initial learning rate of 3e-4, scheduled by WarmupDecayLR with 100 warmup steps. Loss weights are set as $\lambda_{mask}=\lambda_{txt}=1.0$, $\lambda_{policy}=0.1$, $\lambda_{bce}=2.0$, and $\lambda_{dice}=4.0$.
@@ 5.2 Comparison with the State-of-the-Art
Results on RefCOCO(+/g) Dataset. In Table 4, UGround-7B achieves overall stronger performance compared with the recent state-of-the-art POPEN-7B. On RefCOCO, UGround surpasses POPEN-7B by +1.3% (val), +1.5% (testA), and +3.6% (testB), establishing clear new records across all splits. On the RefCOCOg dataset, UGround standout on test with a score of 76.1%, marking a +0.5% improvement over POPEN-7B. These results highlight that UGround consistently outperforms POPEN-7B on most benchmarks.
Results on gRefCOCO Dataset. In Table  5, UGround-7B achieves substantially stronger performance compared with the recent state-of-the-art GSVA-7B (ft) on the gRefCOCO dataset. On the validation set, UGround improvies over GSVA-7B (ft) by +5.99% gIoU, +2.27% cIoU, and +12.10% N-acc, respectively. On Test A, UGround surpasses GSVA-7B (ft) by +3.21% gIoU, +0.94% cIoU, and +8.62% N-acc. These results demonstrate that UGround not only reflects its strength in multi-object segmentation but also excels in correctly identifying empty targets (N-acc).
@@ 5.3 Ablation Study
Table 7: Ablation study on explicit constraints for similarity map $\mathcal{M}$.
@@ 6 Conclusion
Appendix A Supplementary Statements

### v2

@@ 4.1 Policy-Prompted Masking
Algorithm (1) outlines a single forward pass of the PPM. The transformer layers are stacked sequentially. By skipping the subsequent $L$-$\ell^{*}$ layers, the PPM allows layer $\ell^{*}$ to connect directly to SAM in a skip-connection manner, thereby mitigating cumulative errors that arise from layer-by-layer propagation through intermediate correction (i.e., a shortcut that bypasses subsequent layers). Across multiple forward passes (from $1$ to $\mathcal{T}$), the PPM dynamically alters the model’s internal connectivity, analogous to the dropout mechanism: although all layers are virtually connected to SAM, only one path is activated per pass. This stochastic sampling can be interpreted as a form of Monte Carlo uncertainty estimation (Gal and Ghahramani, 2016), which quantifies predictive uncertainty via multiple stochastic forward passes. In the context of PPM, each forward pass activates a different layer-path to SAM, and the ensemble
Algorithm 1 Policy-Prompted Masking Algorithm
[…]
0:  Selected layer $\ell^{*}_{t}$ and similarity map $\mathcal{M}_{t}$
1:  for $t=1,\dots,\mathcal{T}$ do
2:   $s_{\ell}=\bm{h}_{t^{*}}^{(\ell)}\cdot\mathbf{w}_{\ell}$ for $\ell=1,\dots,L$
3:   $\pi_{\theta}(\ell\mid\mathcal{H}_{t^{*}})=\frac{\exp(s_{\ell})}{\sum_{j=1}^{L}\exp(s_{j})}$
4:   $\ell^{*}\sim\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}})$
5:   $\mathcal{S}^{(\ell^{*})}=\mathcal{H}_{\mathbf{z}}^{(\ell^{*})}\,{h_{t^{*}}^{(\ell^{*})}}^{\top}\in\mathbb{R}^{k}$
6:   $\mathcal{M}\in[0,1]^{H\times W}\;\leftarrow\;\mathcal{S}^{(\ell^{*})}$
7:   $M_{\sigma}(u,v)=\iint_{\mathbb{R}^{2}}M(x,y)\,G_{\sigma}(u-x,v-y)\,dx\,dy$
8:   $r=-\big(\mathcal{L}_{bce}(\mathcal{M},M_{\sigma})+\mathcal{L}_{dice}(\mathcal{M},M_{\sigma})\big)$
9:   $b_{t}=\alpha b_{t-1}+(1-\alpha)r$
10:   $\mathcal{L}_{\mathrm{policy}}=-(r-b_{t})\log\pi_{\theta}(\ell^{*}\mid\mathcal{H}_{t^{*}})$
11:   $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}_{\mathrm{policy}}$
12:  end for
13:  return $\{\ell^{*}_{t},\mathcal{M}_{t}\}_{t=1}^{\mathcal{T}}$
of these paths (i.e., 32 in LLaVA) alleviates over-reliance on any single trajectory, enhancing the model’s robustness.
Mask as Prompt. Compared with <SEG> tokens, the similarity map provides more explicit spatial cues, as analyzed in Sec. 3.2. We use the similarity map as a logit mask to prompt SAM. Accordingly, we reformulate Eq. (3) as $\hat{\mathbf{M}}=\;\mathcal{G}_{\mathcal{V}}^{dec}(\mathbf{f},\bm{h}_{seg},\mathcal{M}).$ The similarity map is continuously differentiable, allowing its gradients to be backpropagated through the SAM module for implicit optimization. To further guide the model to focus on target regions, we supervise the similarity map $\mathcal{M}$ by imposing a constraint (e.g., using binary cross-entropy (BCE) or Dice loss) against a soft ground-truth mask $M_{\sigma}$, providing explicit guidance on where the model should “attend”.
@@ 4.2 Training Objectives
5 Experiment
5.1 Experimental Setting
We observe that the center scaling and cropping operation in CLIP may crop out target objects. To address this issue and to align the similarity map with SAM’s input, we adopt SAM’s transformation strategy, scaling the longer side of the image to 336 and padding the image for the CLIP input $\times336$. For results on ReasonSeg, UGround-7B-LLaVA1.5 is trained on 1 NVIDIA A100 GPU 80GB for about 2 days, UGround-13B-LLaVA1.5 for about 13 hours. The training relies solely on ReasonSeg and a referring segmentation dataset ($\sim$10k images). We use LoRA (Hu et al., 2022) for efficient fine-tuning, using lora_r=8lora\_r=8 for 7B and 64 for 13B. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with an initial learning rate of 3e-4, scheduled by WarmupDecayLR with 100 warmup steps. Loss weights are set as $\lambda_{mask}\!=\!\lambda_{txt}\!=\!1.0$, $\lambda_{policy}\!=\!0.1$, $\lambda_{bce}\!=\!2.0$, and $\lambda_{dice}\!=\!4.0$.
@@ 5.2 Comparison with the State-of-the-Art
Results on RefCOCO(+/g) Dataset. In Table 4, UGround-7B achieves stronger overall performance than the recent state-of-the-art GLaMM-7B. On RefCOCO, UGround surpasses GLaMM-7B by +1.1% on val and +0.8% on testB. On RefCOCO+, it exceeds GLaMM-7B by +1.0% on testB and remains competitive on val and testA. On RefCOCOg, UGround reaches 76.1% on test(U), improving +1.2% over GLaMM-7B, highlighting its generalization across classical referring expression segmentation benchmarks.
Results on gRefCOCO Dataset. In Table  5, UGround-7B achieves substantially stronger performance compared with the recent state-of-the-art GSVA-7B (ft) on the gRefCOCO dataset. On the validation set, UGround improves over GSVA-7B (ft) by +5.99% gIoU, +2.27% cIoU, and +12.10% N-acc, respectively. On Test A, UGround surpasses GSVA-7B (ft) by +3.21% gIoU, +0.94% cIoU, and +8.62% N-acc. These results demonstrate that UGround not only reflects its strength in multi-object segmentation but also excels in correctly identifying empty targets (N-acc).
@@ 5.3 Ablation Study
Table 7: Ablation study on loss constraints for similarity map $\mathcal{M}$.
@@ 6 Conclusion
Impact Statement
This paper presents UGround to advance a unified paradigm for visual grounding. By dynamically selecting intermediate layers in unrolled transformers and using “mask as prompt” to provide explicit spatial cues to a segmentation model (e.g., SAM), UGround supports a broad range of settings within one system, spanning traditional referring expression segmentation to reasoning segmentation, single-target to multi-target queries, and rejection under false-premise (empty-target). In terms of societal impact, visual grounding can benefit human–AI interaction for image understanding, data annotation, and content editing. However, they may also be misused for privacy-invasive or surveillance applications (e.g., automatically localizing and extracting sensitive regions in personal imagery), or to facilitate the creation and dissemination of misleading content (e.g., precise cutouts and compositing of specific subjects). To mitigate these risks, we adhere to established community standards for responsible deployment. In particular, we encourage auditability in real-world applications and transparency in model design to foster trust.
Appendix A Notations
The notations used throughout this paper are summarized in Table 9 for clarity.
Table 9: Key notations used in this paper.
Notation Description
$\mathcal{G}_{\theta}$ LMMs-based segmentation model, including a multi-modal LLM $\mathcal{G}_{\mathcal{T}}$ and a visual backbone model $\mathcal{G}_{\mathcal{V}}$, $\mathcal{G}_{\theta}=\mathcal{G}_{\mathcal{V}}\circ\mathcal{G}_{\mathcal{T}}$
$\mathcal{H}^{(\ell)}$ the hidden states at the $\ell$-th transformer layer
$\bm{h}_{t^{*}}^{(\ell)}$ the hidden state of <SEG> token at layer $\ell$-th (position $t^{*})$
$\hat{\mathbf{M}}$ the predicted binary segmentation mask
$M$ the ground-truth mask
$M_{\sigma}$ the softened label obtained by applying a Gaussian kernel to $M$
$\varphi(\cdot)$ a multilayer perceptron (MLP) projection layer
$\mathcal{M}$prompt mask as a prompt for SAM decoder $\mathcal{G}_{\mathcal{V}}^{dec}$
<SEG>prompt <SEG> token as prompt for SAM decoder $\mathcal{G}_{\mathcal{V}}^{dec}$
$\mathcal{M}$IoU the overlap between the similarity map and the ground-truth mask
$\mathcal{S}^{(\ell)}$ the similarity scores at the layer $\ell$-th,
$\mathcal{M}$ the soft logit mask or similarity map $\mathcal{M}\in[0,1]^{H\times W}$
$\pi_{\theta}(\cdot\mid\mathcal{H}_{t^{*}})$ the policy distribution conditioned on $\mathcal{H}_{t^{*}}$
$G_{\sigma}(x,y)$ a 2D Gaussian kernel with $\sigma$, $G_{\sigma}(x,y)\propto\exp(-\frac{x^{2}+y^{2}}{2\sigma^{2}})$
Appendix B Analysis
Complexity Analysis. To justify the complexity of RL-based layer selection (RL-gating), we compare it against a simpler non-RL alternative—soft-gating, i.e., a learned soft-attention mechanism that produces a weighted average over layer outputs. We report results on ReasonSeg val set. In Table 10, RL-gating consistently outperforms soft-gating, with gains of +1.47% in gIoU and +1.34% in cIoU.
Table 10: RL-gating vs. soft-gating.
Exp. ID Method gIoU cIoU
1 soft-gating 64.66 70.73
2 RL-gating 66.13 72.07
Runtime Analysis. To assess the computational costs of UGround, we report training and inference latency on a single NVIDIA A100-SXM4-40GB GPU in Table 11 and Table 12. We train on the ReasonSeg train split over 400 samples for 1 epoch (200 steps), with a batch size of 2. Training Latency (s) refers to the average time per iteration during the 200-step training process (including forward, backward, and optimizer step), and Memory Usage refers to the MaxMemAllocated observed within the 200 steps. We also report the model’s Trainable Percentage (Trainable Parameters / Total Parameters). For inference, we evaluate on the ReasonSeg test split over 799 samples, with a batch size of 1. Speed (FPS) refers to the average frames per second computed over the 799 samples. In Table  11, UGround exhibits a higher training latency (2.97s) than LISA (Lai et al., 2024)(1.26s) due to the explicit optimization of the similarity map during the backward pass. In Table  12, UGround achieves a comparable inference runtime to LISA (Lai et al., 2024) and SESAME (Wu et al., 2024).
Table 11: Comparing the training cost of our UGround to state-of-the-art methods.
Model Training Latency (s) Memory Usage (GB) Trainable (%) Trainable Params Total Params
SESAME (Wu et al., 2024) 1.11 23.15 3.73% 288.25M 7.73B
PixelLM  (Ren et al., 2024) 1.22 23.02 5.25% 375.72M 7.16B
LISA (Lai et al., 2024) 1.26 23.68 3.74% 288.25M 7.71B
GSVA (Xia et al., 2024) 2.13 25.73 3.73% 288.26M 7.73B
UGround 2.97 28.16 3.72% 288.40M 7.75B
Table 12: Comparing the runtime speed of our UGround to state-of-the-art methods.
Model GSVA (Xia et al., 2024) SESAME (Wu et al., 2024) LISA (Lai et al., 2024) PixelLM (Ren et al., 2024) UGround
Speed (FPS) 3.98 4.64 4.68 9.24 4.12
Generalization Analysis. To demonstrate broader generalization of UGround, we report results on MUSE benchmark. MUSE aims for multi-target reasoning segmentation (Ren et al., 2024). In Table 13, UGround-7B consistently outperforms PixelLM-7B across all splits of the MUSE benchmark. It improves the Val overall cIoU by +5.7% and boosts the challenging “few-targets” cIoU by +10.5%. On the Test overall set, UGround also achieves a +4.1% cIoU gain, yielding a clear performance margin over PixelLM-7B. Also, we report results on augmented FP-Refcoco(+/g) under false premises. In Table 14, UGround outperforms READ across all FP-Refcoco(+/g) splits, with gains of +1.59% to +2.99% in “see” accuracy and +1.30% to +2.43% in segment scores (cIoU).
Table 13: Comparison on MUSE benchmark. Numbers are cited from PixelLM (Ren et al., 2024).
Method w/o SAM Val Test
overall few targets many targets overall
gIoU cIoU gIoU cIoU gIoU cIoU gIoU cIoU
SEEM (Zou et al., 2023b) 13.6 16.2 23.6 24.9 8.5 13.2 11.7 15.7
LISA-7B (Lai et al., 2024) 18.8 29.0 24.7 36.5 9.6 24.5 12.8 27.1
[…]
PixelLM-7B (Ren et al., 2024) 42.6 50.7 44.6 59.2 37.7 42.8 39.2 46.3
UGround-7B 52.4 56.4 46.2 69.7 38.8 46.9 40.1 50.4
Table 14: Comparisons of the state-of-the-art “see” and “segment” results on augmented FP-Refcoco(+/g) val set. “FP”(False Premise) denotes a query for an object absent from the image.
Method FP-RefCOCO FP-RefCOCO+ FP-RefCOCOg
See Segment See Segment See Segment
LISA-7B (Lai et al., 2024) 51.36 44.00 51.32 39.62 51.25 39.64
Cascading (Wu et al., 2024) 75.59 55.18 75.03 48.64 76.07 49.98
SESAME (Wu et al., 2024) 79.84 57.93 80.00 50.81 81.78 53.79
READ (Qian et al., 2025) 82.87 61.50 83.51 54.54 84.67 56.12
UGround 85.86 62.80 85.10 56.03 86.86 58.55
Appendix C Discussions
C.1 Justification for Reinforcement Learning (RL)
Reinforcement Learning (RL) is introduced to address the hard-boundary problem in layer selection, enabling the model to explicitly choose a layer (e.g., layer 26) in a soft, differentiable manner during training. Because hard-boundary layer selection is non-differentiable, using RL is necessary.
An intuitive approach is to train a classifier for dynamic layer selection using cross-entropy loss, followed by soft-gating via a weighted average over all layer outputs. We call this simpler alternative CE + soft-gating. Here, the layer-selection label is assigned based on the overlap (IoU) between each layer’s similarity map (layers 0–31) and the ground-truth mask. The classifier is trained, and its output scores serve as soft gates to weight and fuse the transformer layers (0–31) into a unified representation, which is then fed into SAM for mask decoding. However, CE + soft-gating suffers from the following issues:
1) CE + soft-gating is numerically suboptimal. Consider a scenario after training for 1–2 epochs where the classifier predicts the following scores for layers 0–31: ($0.01,0.004,\ldots,0.7,0.1,0.1]$). Since 0.7 is the maximum score, the 29-th layer is identified as the best layer. However, if we use the full score vector ($0.01,0.004,\ldots,0.7,0.1,0.1]$) as soft-gating weights to fuse the embeddings of layers 0–31, the best embedding from layer 29 will inevitably be mixed with features from all other layers (0–28 and 30–31), making it no longer optimal. Soft-gating inherently prevents the remaining layers from receiving zero weight. In fact, as shown in Fig. 4 ( and also discussed in (Zhang et al., 2025)), transformer layers are activated only at a few specific depths. Assigning non-zero weights to inactive layers dilutes the contribution of these key activated layers. Therefore, a hard boundary is required—that is, explicitly selecting a single layer (e.g., assigning 100% weight to layer 29) rather than distributing fractional weights (e.g., 60% on layer 29). This is inherently a discrete decision problem, for which reinforcement learning (RL) provides a principled and trainable solution.
2) The training objective of the layer-selection classifier (i.e., CE loss) does not align with the ultimate goal of optimizing the gates through backpropagation from the downstream SAM mask decoding (i.e., use soft-gating without an explicit loss on the gates). The layer-selection classifier explicitly trains the gate to predict which layer is the most important, whereas using soft weights to fuse all layers aims to optimize the averaged representation, without explicitly enforcing the gate to select the optimal layer. Besides, soft-gating presumably ignores inter-layer dependencies in the transformer. By weighting and summing all layers, it actually performs a soft ensemble, disrupting the original layer-by-layer structure of the transformer.
3) Computational cost: CE + soft gating vs. RL-gating. Given $\bm{h}_{seg}\in\mathbb{R}^{4096}$ and $\bm{h}_{img}\in\mathbb{R}^{576\times 4096}$, the similarity map is computed as: $\bm{h}_{seg}\times\bm{h}_{img}^{T}=(1\times 4096)\times(576\times 4096)^{T}=576$, followed by reshaping into $24\times 24$, and then interpolating to $256\times 256$ to obtain $\mathcal{M}$. If using the CE + soft-gating approach, one must compute the similarity map for all $32$ layers of LLaVA, resulting in $32\times 256\times 256$ maps, and then compute the overlap (IoU) with the ground-truth mask as the supervision label. In contrast, with RL, we compute $\bm{h}_{seg}\times\bm{w}^{T}=(1\times 4096)\times(32\times 4096)^{T}=32$, yielding a 32-dimensional similarity logits vector, followed by REINFORCE to select the layer index, e.g., ($\ell^{*}=28$). $\bm{w}^{T}\in\mathbb{R}^{32\times 4096}$ denotes the layer-gating weights, which are the only learnable parameters in UGround. The reward $r$ only requires computing the similarity map for a single layer.
C.2 Costs of Mask as Prompt
Compared to the <SEG> token, Mask as Prompt (MasP) is more effective, as it provides explicit spatial cues via its activation regions. Here, we analyze the computational costs associated with Mask as Prompt. We use SAM (Kirillov et al., 2023) for mask decoding. The original paper describes its prompt encoder as follows:
“Prompt encoder. We consider two sets of prompts: sparse (points, boxes, text) and dense (masks). We represent points and boxes by positional encodings summed with learned embeddings for each prompt type, and free-form text with an off-the-shelf text encoder from CLIP (Radford et al., 2021). Dense prompts (e.g., masks) are embedded using convolutions and summed element-wise with the image embedding.”
Table 15: Mask prompt dimension changes through convolutional layers in SAM.
Layer Output shape
Input mask: $256\times 256$ (B, 1, 256, 256)
Conv2d(1 $\rightarrow$ 4, stride=2) (B, 4, 128, 128)
Conv2d(4 $\rightarrow$ 16, stride=2) (B, 16, 64, 64)
Conv2d(16 $\rightarrow$ 256, stride=1) (B, 256, 64, 64)
That is, SAM accepts a mask prompt of size $256\times 256$, which is transformed by the convolutional layers in Table 15. Thus, the final mask embedding has dimensions [B, 256, 64, 64]. In SAM’s implementation, even when the mask-as-prompt input is not provided, i.e., when only the text <SEG> prompt is used (mask=None), SAM still generates a set of learnable embeddings of the same shape as a placeholder, serving as the mask-embedding representation. Therefore, using Mask as Prompt does not introduce any additional parameter overhead.
C.3 Analogous to Monte Carlo Dropout
Analogous to Monte Carlo Dropout. Consider the query: “Can you segment the apple, orange, and banana in the picture ?” UGround responds: “Yes, apple, orange, and banana are <SEG>, <SEG>, <SEG>.” That is, there are three objects to segment in a single image. During one forward pass, our SSC module (RL) selects a layer from 0–31 for each of $\texttt{<SEG>}_{apple}$, $\texttt{<SEG>}_{orange}$, and $\texttt{<SEG>}_{banana}$ as input to SAM. Suppose layers (15, 28, 19) are selected, that is, $\texttt{<SEG>}_{apple}^{(15)}\rightarrow$ SAM, $\texttt{<SEG>}_{orange}^{(28)}\rightarrow$ SAM, $\texttt{<SEG>}_{banana}^{(19)}\rightarrow$ SAM. In the ($\mathcal{T}$-1)-th forward pass, the selections are $\texttt{<SEG>}_{apple}^{(16)}\rightarrow$ SAM, $\texttt{<SEG>}_{orange}^{(20)}\rightarrow$ SAM, $\texttt{<SEG>}_{banana}^{(14)}\rightarrow$ SAM. In the $\mathcal{T}$-th forward pass, the selections are $\texttt{<SEG>}_{apple}^{(13)}\rightarrow$ SAM, $\texttt{<SEG>}_{orange}^{(17)}\rightarrow$ SAM, $\texttt{<SEG>}_{banana}^{(26)}\rightarrow$ SAM.
During the $\mathcal{T}$ forward passes, $\texttt{<SEG>}_{apple}$ from layers 15, 16, and 13 are independently connected to SAM, and the SSC module dynamically alters the model’s internal connectivity. This can be analogized to the dropout mechanism in neural networks: All layers are virtually connected to SAM, but only one path is activated per forward pass. Since the SSC module performs stochastic sampling, this can be interpreted as a Monte Carlo uncertainty estimation (Gal and Ghahramani, 2016). Unlike soft-gating, the SSC module enables each layer to connect independently to SAM, resulting in 32 independent subnetworks in an ensemble-style design. As only one layer is activated at a time, the computational overhead is comparable to selecting a fixed last layer, thus avoiding expensive computational costs.
Appendix D Qualitative Results
Layer-wise visualization of the Transformer across layers 1-40. Fig. 46 present a qualitative analysis of similarity map activations across layers 1–40, illustrating their alignment with the ground-truth mask. Each figure contrasts heatmaps generated by prior works using a fixed last hidden layer (a) with those from our proposed Policy-Prompted Masking method (b). Unlike the baseline approach (a), our method introduces stochastic layer selection: in each forward pass, a layer $\ell^{*}$ is randomly chosen, and multiple $\mathcal{T}$ stochastic samples result in an approximate Bayesian uncertainty estimation, akin to Monte Carlo Dropout. This sampling mechanism explores diverse pathways, alleviates over-reliance on any single trajectory, and virtually forms an ensemble of subnetworks, thereby enhancing robustness. As a result, the similarity maps in (b) show activations that are sharply concentrated on the target object in intermediate and later layers, yielding more accurate alignment with the ground-truth mask compared to the diffuse and noise-prone activations observed in (a).
Visualization of UGround on benchmarks. Fig. 78 show qualitative results on ReasonSeg test set. Fig. 8 demonstrates UGround’s capability in semantic segmentation, i.e., assigning pixel-level labels for a target semantic category (with background as 0). Fig. 910 show qualitative results on gRefCOCO (Liu et al., 2023) test sets A and B, which demonstrate UGround’s capability in instance segmentation, i.e., predicting instance-level masks for individual object instances (with separate masks for objects of the same category). Fig. 11 shows qualitative results on FP-RefCOCO(+/g) val set, which demonstrates UGround’s capability in handling false-premise queries, i.e., queries that refer to objects not present in the given image (empty target). Notably, UGround is capable of correcting false premises, i.e., directly refusing queries when the target object is absent and optionally suggesting look-alike objects (Wu et al., 2024) as shown in Fig. 11.
Appendix E Implementation Details
The UGround code repository. Currently, UGround supports 8 dataset types, namely, A: sem_seg, B: refer_seg, C: neg_refer_seg, D: correct_refer_seg, E: vqa, F: reason_seg, G: reason_seg_plus, and H: multi_reason_seg. In particular,
A: sem_seg: ade20k $||$ cocostuff $||$ pascal_part $||$ paco_lvis $||$ mapillary
B: refer_seg: refclef $||$ refcoco $||$ refcoco+ $||$ refcocog $||$ refzom $||$ grefcoco
C: neg_refer_seg: R-Refcoco $||$ R-Refcoco+ $||$ R-Refcocog
D: correct_refer_seg: fprefcoco $||$ fprefcoco+ $||$ fprefcocog
E: vqa: llava_instruct_150k
F: reason_seg: ReasonSeg||train
G: reason_seg_plus(LISA++): instance_seg $||$ cot $||$ conversations $||$ caption
H: multi_reason_seg(muse): MultiReasonSeg||train
Our UGround is trained in a mixed fashion on these 8 types of “in-the-wild” datasets as needed. As shown in Fig. 13, the UGround project is organized with a model directory for core implementations, a scripts directory containing training and testing scripts, and a configs directory for configuration files. The main model architecture is defined in UGround.py within the UGroundForCausalLM class, which extends LlavaLlamaForCausalLM and integrates a Segment Anything Model (SAM) for mask generation. A central component of this architecture is the PolicyPromptedMasking class, detailed in PPM.py, which is instantiated and used by the UGroundForCausalLM class to dynamically select hidden state features from different network layers. The selection policy is trained using the REINFORCE algorithm with either an EMA or a critic-based baseline. All codes and models are publicly available at https://github.com/rui-qian/UGround.
Appendix F Ethics and Societal Impact